OpenStudy (anonymous):

The sum of three times a number and 12 is equal to half of the same number plus 22. What is the number?

OpenStudy (anonymous):

number:x so 3x + 12 = (x/2) + 22; multiply both sides by 2 and you get 6x + 24 = x +44 and solve for x and you get 5x=20 so x=4